Question 2 1 pts Assuming that the price of labor is w -10 and the price of capital is k -40, use Table 1 from the Excel sheet labeled Exercise2Table.xlsx to do the following: 1. Complete Table 3 2. If the firm can sell the good being produced at $73 per unit, determine whether the firm should produce and sell 1, 2 or 4 units. 3. Complete Table 4 by determining the cheapest way to produce the given amounts of output in Table 4 (using the information you have from Table 1)

TABLE 1: Data 1.00 4.00 2.00 1.00 1.56 1.83 2.05 2.25 2.43 2.60 2.76 4.00 2.00 8.00 9.00 16.00 8.00 9.00 16.00 9.00 16.00 8.00 16.00 16.00 36 16 0 32 36 100 18 16 64 12.5 16 16 25 25 32 49 64 36 0

TABLE 3: Short Run Costs FC VC TC MC AFC AVC ATC 1.00 1.56 1.83 2.05 2.25 2.43 2.60 2.76 4.00 5.83 9.00 4 4 4 4 4 16 32 64

TABLE 4: Long Run Costs TC MC ATC 0.00 1.00 2.00 3.00 4.00 8.00 9.00 16.00
Answer #1

Consider the given problem here we have given all the information in the given question. So, here “K” is the fixed input and “L” is the variable input. So, the “FC” is given by “R*K=40*K” and the “VC” is given by “W*L = 10*L”, => the total cost is given by the sum of “FC” and “VC”.

FC-R*k VC-WL TC-FC+VC MC-TC(n)-TC(n-1)/(Qn-Qn-1) AFC-FC/Q AVC-VC/Q ATC AFC+AVO 160 160 160 160 160 160 160 160 160 160 160 160 170 160 1.00 1.56 1.83 2.05 2.25 2.43 2.60 2.76 4.00 5.83 9.00 160 108.97 98.36 92.68 88.89 86.42 84.62 83.33 1 10 17.86 102.56 37.04 45.45 50.00 55.56 58.82 62.50 72.58 87.43 6.41 10.93 14.63 17.78 20.58 23.08 25.36 2 87.43 78.05 4 210 65.84 61.54 57.97 4 230 16 160 320 320 27.44 17.78 32 54.89 82.33 100.95 88.89

Now, the “MC” is the additional cost to produce additional unit of output. So, Now, once we get “FC”, “VC” and “TC” we can easily calculate “AFC”, “AVC” and “ATC”.

Now, if “Q=1”, => “TC=160” and the corresponding revenue is “R=P*Q=73*1=73”, => the profit is “R-TC=73-160 = (-87)”. Less than the “FC”, => the firm will continue its production. Now, if “Q=2.05”, => “TC=190” and the corresponding revenue is “R=P*Q=73*2.05 = 149.65”, => the profit is “R-TC = 149.65 -190 = (-40.35)” less than the “FC”, => the firm will continue its production. Similarly, for “Q = 4”, => “TC = 320” and the corresponding revenue is “R=P*Q=73*4 = 292”, => the profit is “R – TC = 292 - 320 = (-28)” less than the “FC”, => the firm will continue its production.

806678 142111222 2.5 76 4 2 8 9 16 8 , 16 , 16 8 16 16 66428968969686 饒 468012345676026086 33011 64 46040 112444444444889956655294 1

饒 TC MC ATC 160 160 160 160 C) 320 80 80 80 720 16 1280

Now, from the 1st table we can find out the associated cost for all input combinations and choose the one having lowest cost. Consider the following table shows the LR cost and the “MC” and “ATC” for each units of output produced.